• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 10:42:04 on July30,2003

    Entry number 108172

    This entry is a followup to: 107914

    keyword=2nd correction to owl shift run summary 7/29/03

    runs 2974/22974 were started stopped @ 0209:0222 and had 3M/206K events.
    it was a Pol 3He, 0 deg run. run 22974 had a really high deadtime > 60%.

    run 2975 failed at the run end. run 22975 had a 100% deadtime. this was the set of runs when the DAQ problem really started.